OpenStudy (anonymous):

Given A (-6,4) and M(-1, -1) find B. 1) (6,4) 2) (-6, -4) 3) (-4, 6) 4) (4, -6)

OpenStudy (tad1):

Use mid-point formula. (-6 +x)/2= -1 and (4 + y)/2 + -1. Solve for x and y.

OpenStudy (anonymous):

Find the distance to the midpoint, then just go twice that value to get to the other end. For x: Mx-Ax= (-1-(-6))=5, then just add 5 to the midpoint, so -1+5=4 for y: My-Ay=(-1-(4))=-5, then just add -5 to the midpoint, so -1-5=-6 Result is (4,-6)

OpenStudy (tad1):

That would give you the length, but not the coordinates of the point.